# Pretrained Hybrids with MAD Skills

### v1

@@ Abstract
While Transformers underpin modern large language models (LMs), there is a growing list of alternative architectures with new capabilities, promises, and tradeoffs. This makes choosing the right LM architecture challenging. Recently-proposed hybrid architectures seek a best-of-all-worlds approach that reaps the benefits of all architectures. Hybrid design is difficult for two reasons: it requires manual expert-driven search, and new hybrids must be trained from scratch. We propose Manticore, [^2] The Manticore is a fearsome human/lion/scorpion hybrid from Persian mythology. a framework that addresses these challenges. Manticore automates the design of hybrid architectures while reusing pretrained models to create pretrained hybrids. Our approach augments ideas from differentiable Neural Architecture Search (NAS) by incorporating simple projectors that translate features between pretrained blocks from different architectures. We then fine-tune hybrids that combine pretrained models from different architecture families—such as the GPT series and Mamba—end-to-end. With Manticore, we enable LM selection without training multiple models, the construction of pretrained hybrids from existing pretrained models, and the ability to program pretrained hybrids to have certain capabilities. Manticore hybrids outperform existing manually-designed hybrids, achieve strong performance on Long Range Arena (LRA) tasks, and can improve on pretrained transformers and state space models.
@@ 1 Introduction
Transformers are the workhorse architecture for large language models and beyond, powering a vast collection of foundation models. While for years it appeared that the Transformers family would remain the undisputed standard, a recent Cambrian explosion of proposed architectures has taken place. Many of the new architectures achieve subquadratic complexity—in contrast to the quadratic complexity of self-attention in Transformers—by using local or linear attention [9, 6, 3, 45], or resurrecting and scaling recurrent networks [6, 9, 31], or by building on state-space modeling principles [13, 33, 32, 11, 14]. These approaches potentially promise to overturn the dominance of Transformers through more efficient training and inference.
However, no single new model is a clear overall winner when varying data modalities, tasks, and model sizes. Comparing architectures on a fixed task is fraught with difficulties [2]. Even if these issues are overcome, practitioners would have to experiment with and evaluate every architecture for each new task—an expensive proposition. Instead, seeking a best-of-all-worlds approach, researchers have proposed the use of hybrid models that mix multiple architectures. These hybrids, such as the MambaFormer [29]—a mix of the popular Mamba with a Transformer—have shown potential in maintaining the desirable properties of multiple model classes.
@@
  • Failure to Use Pretrained Models. It is unclear how to integrate pretrained model components from models with different architectures. Pretrained models are a key advantage of foundation models. However, due to compatibility issues, hybrids are often trained from scratch, leading practitioners to resort to small hybrids in limited settings or incur high costs.
A potential solution to the latter challenge is the use of model merging techniques [42, 44, 41, 15, 8, 16], some of which can operate cross-architecture [1, 12]. Unfortunately, such tools are embryonic–they are expensive and it is unclear how well they work with the diverse types of architectures a user may seek to build a hybrid from.
We propose a framework for automatically designing hybrid architectures that overcomes these obstacles. Our approach is inspired by principles from neural architecture search (NAS), but applies these at the level of LM blocks rather than convolutional cells [23, 21] or operations [36, 35]. The resulting framework is simple and tractable. It sidesteps merging different architectures by using simple linear projectors to translate between the “languages” spoken by various architectures. This enables us to include blocks from many different architectures with little to no changes required. In addition, inspired by the mechanistic architecture design (MAD) framework [34], we show how our framework can learn hybrid architectures via MAD that transfer to new tasks.
Concretely, our proposed system, Manticore:
@@
    Automatically selects language models, without training several models from scratch,
@@
    Automatically constructs pretrained hybrids without evaluating the entire search space,
@@
    Provides a technique for programming hybrids to have certain skills without full training.
Experimentally, our automatically designed hybrids compete with existing hybrids and models on the MAD tasks [34] and Long Range Arena (LRA) [38], we produce pretrained hybrids that can improve downstream fine-tuning performance, and we can program hybrids using the MAD tasks.
@@ 2 Related work
Language Model Architectures: Transfomers and Beyond. Transformers are currently the dominant LM architecture. The success of the “vanilla” architecture introduced by Vaswani et. al. [40] has led to many proposed variations. The quadratic complexity of the base self-attention operation has inspired the search for alternative architectures that offer comparable performance with subquadratic complexity. One line of work builds off state-space models, with variations made to enable language modeling [32, 33, 13, 3]. Another line of work involves linear attention by formulating transformers as RNNs and expressing self-attention as a kernel dot-product [17]. Other new approaches increase the expressivity of this formulation with data-dependent gating [43].
Our work does not propose a new architecture. Instead, we focus on the idea that practitioners should be able to take advantage of new architectures in a transparent way.
Neural Architecture Search & Mechanistic Search. Neural architecture search (NAS) techniques are used to automatically search for optimal architectures. These techniques have produced state-of-the-art models in several domains. Much of the challenge in NAS is the complexity of search; in the most standard form, NAS involves a difficult bilevel optimization over a large search space. Much effort has been aimed at reducing costs, often via continuous relaxations of search spaces, with techniques like DARTS [23] and DASH [36].
Using NAS to discover architectures for language modeling—and especially those that may rival Transformers—has thus far been hard. A promising approach is the MAD framework [34] , which uses “mechanistic tasks” (synthetic tasks organized around simple principles) to search for high-quality architectures. While we do not seek to discover new architectures, we are inspired by this approach in our effort to search for hybrid architectures.
@@ 3 Methods
We now describe Manticore, our framework for automatically designing hybrid architectures by mixing components of pretrained models. Manticore relies on projectors to align features across architectures, then applies a convex combination to the aligned features, as summarized in Figure 1. In Section 3.1, we discuss and formally define the structure of Manticore hybrids, including the projectors and convex combination mixture weights, as well as how both of these components are used within Manticore. In Section 3.2, we detail the search procedures (inspired by NAS) and training routines involved in pretraining, fine-tuning, and programming hybrids.
Figure 1: Our proposed Manticore framework, which enables: (1) cross-architecture LM selection, (2) the construction of pretrained hybrids, and (3) the ability to program hybrids to have certain skills.
@@ 3.1 The Structure of Manticore Hybrids
3.1.1 Component Models
We refer to a model that is used in Manticore as a component model. Any modern LM can be used as a component model in our framework. In this section, we will formally define the general high-level structure of the component models that we support. For an LM $M$ with model embedding dimension $d_{M}$ on a sequence of $t$ tokens from a set $\mathcal{V}$, denoted $x=(x_{1},...,x_{t})\in\mathcal{V}^{t}$, a forward pass $M(x)$ is typically computed using the following recipe:
@@ 3.1.1 Component Models
This recipe applies to virtually all modern transformer-based LMs, recurrent models, and state-space models. Our framework supports all of these, and any other architecture that follows this recipe.
3.1.2 Projectors
Suppose we have two pretrained component models, $M$ and $M^{\prime}$. In general, even if the model dimensions are the same for both models ($d_{M}=d_{M^{\prime}}$), blocks from $M$ and $M^{\prime}$ may not be directly compatible, as their input and output features are likely to be very different. It is also possible that $d_{M}\neq d_{M^{\prime}}$, in which case composing blocks from $M$ and $M^{\prime}$ is not even well-defined.
Figure 2: Examples of component models used in Manticore. Transformer and Mamba component models are shown.
Figure 3: The projectors with residual connections used in Manticore, used for translating features between pretrained blocks of different component models.
To overcome this issue, we apply projectors to both the inputs and the outputs of a block (or a sequence of blocks, discussed in Section 3.1.4) that we wish to combine in Manticore hybrids. Overall, our goal in designing projectors is to enable the blocks of $M$ and $M^{\prime}$ to speak a common language, such that their features are compatible and can be reused in the resulting hybrid model. This is conceivably challenging—the mapping between feature spaces could be highly nonlinear and might require a lot of task-specific data to adequately learn the mapping. So do projectors need to be heavyweight, data-hungry, highly nonlinear objects? Fortunately, the answer is no—we find that a simple linear transformation with a gated residual, pretrained on general language data, is sufficient.
@@ 3.1.2 Projectors
[…]
$\circ M_{(k)\text{Block}}^{(\ell)}\circ\text{Proj-in}^{(\ell)}_{(k)}\bigg{)}:\mathbb{R}^{t\times d_{\text{max}}}\to\mathbb{R}^{t\times d_{\text{max}}}.$
@@
For input $x\in\mathbb{R}^{t\times d_{M_{(k)}}}$ we parameterize each projector as a linear transformation with gated residual:
[…]
Respectively, $\text{Trunc}(\cdot;d)$ and $\text{Pad}(\cdot;d)$ truncate and zero-pad input to dimension $d$, and $\text{Linear}_{d\to d^{\prime}}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ is a learnable linear transformation. Gating weights are parameterized as $\alpha\in[0,1]$.
@@
In total, where $\alpha\in\Delta^{K-1}$ and $I_{k}$ is a length-$n_{k}$ vector of block indices from component model $k$, we define the output of the block sequence defined by $I_{k}$ as
$h_{k}(x;\alpha_{k},I_{k})=\left(\text{Proj-out}^{(I_{k,n_{k}})}_{(k)}\circ M_{(k)\text{Block}}^{(I_{k,n_{k}})}\circ...\circ M_{(k)\text{Block}}^{(I_{k,1})}\circ\text{Proj-in}^{(I_{k,1})}_{(k)}\right)(x;\alpha_{k}).$
3.1.3 Mixture Weights
@@
Next, we would like to mix the activations of different component models’ block sequences, in a way that allows us to learn how much influence the blocks from each component model will have on the overall hybrid model. Learning the amount of influence that each block sequence should have on the overall hybrid is critical—if certain blocks produce less helpful features, we need a way to down-weight them. Conversely, we want to use the best blocks in our hybrid as much as possible—we want to up-weight these helpful blocks. Overall, a parameterization that allows us to learn these weights should lead to better hybrids.
We do this by taking a convex combination of the projectors’ outputs: given the projected features $h_{k}(x;\alpha_{k},I_{k})$ for each component model $k\in[K]$, we output a convex combination of projected features
$\text{Mix}_{\alpha}(x;I_{1},...,I_{K})=\sum_{k\in[K]}\alpha_{k}h_{k}(x;\alpha_{k},I_{k}).$
Figure 4: Mixture weights used in Manticore, which learn how much influence component model blocks should have.
@@
We reuse the convex combination weights as the gating weights in the projectors. This choice yields the convenient property that when the mixture weights $\alpha$ are set to one in index $k$ and zero everywhere else, the Mix function exactly computes a sequence of blocks from component model $k$ while ignoring the projectors and the blocks from other component models. We adopt a popular parameterization for mixture weights from the NAS literature [23]: we parameterize $\alpha$ using a softmax of scalars. That is, we define $\alpha_{k}:=\frac{\exp(a_{k})}{\sum_{j\in[K]}\exp(a_{j})}$ for all $k\in[K]$.
3.1.4 Manticore
@@
We are now ready to define our overall hybrid architecture. We seek to create a hybrid from $K$ component models, $M_{(1)},...,M_{(K)}$, each with a potentially different number of blocks, denoted $L_{M_{(k)}}$ for component model $k$. We fix $L$ to be the number of Manticore blocks, where $L$ is a common factor of each of the depths $L_{M_{(k)}}$, for all $k\in[K]$—we treat this choice of factor as a hyperparameter. For each of the $L$ Manticore blocks, we want to mix a sequence of blocks from each of the $K$ component models. We also want the number of blocks from each model $k\in[K]$ that are allocated to a single Manticore block to be evenly spread out throughout the $L$ Manticore blocks—this is why we require $L$ to be a factor of $L_{M_{(k)}}$.
For each component model $k\in[K]$, divide the indices of the blocks $[L_{M_{(k)}}]$ evenly into $L$ contiguous parts, denoted as $[L_{M_{(k)}}]=(I_{k,1},...,I_{k,L})$. Then, adopting the notation from our component models, a Manticore block is defined as

### v2

@@ Abstract
While Transformers underpin modern large language models (LMs), there is a growing list of alternative architectures with new capabilities, promises, and tradeoffs. This makes choosing the right LM architecture challenging. Recently proposed hybrid architectures seek a best-of-all-worlds approach that reaps the benefits of all architectures. Hybrid design is difficult for two reasons: it requires manual expert-driven search, and new hybrids must be trained from scratch. We propose Manticore, [^2] The Manticore is a fearsome human/lion/scorpion hybrid from Persian mythology. a framework that addresses these challenges by automating the design of hybrid architectures while reusing pretrained models to create pretrained hybrids. Our approach augments ideas from differentiable Neural Architecture Search (NAS) by incorporating simple projectors that translate features between pretrained blocks from different architectures. We then fine-tune hybrids that combine pretrained models from different architecture families—such as the GPT series and Mamba—end-to-end. With Manticore, we enable LM selection without training multiple models, the construction of pretrained hybrids from existing pretrained models, and the ability to program pretrained hybrids to have certain capabilities. Manticore hybrids match existing manually designed hybrids, achieve strong performance on Long Range Arena, and improve on pretrained transformers and state space models on various natural language tasks.
@@ 1 Introduction
Transformers are the workhorse architecture for large language models and beyond, powering a vast collection of foundation models. While for years it appeared that the Transformers family would remain the undisputed standard, a recent Cambrian explosion of proposed architectures has taken place. Many of the new architectures achieve subquadratic complexity—in contrast to the quadratic complexity of self-attention in Transformers—by using local or linear attention [9, 6, 3, 48], resurrecting and scaling recurrent networks [6, 9, 34], or by building on state-space modeling principles [14, 36, 35, 11, 15]. These approaches potentially promise to overturn the dominance of Transformers through more efficient training and inference.
However, no single new model is a clear overall winner when varying data modalities, tasks, and model sizes. Comparing architectures on a fixed task is fraught with difficulties [2]. Even if these are overcome, practitioners would have to experiment with and evaluate every architecture for each new task—an expensive proposition. Instead, seeking a best-of-all-worlds approach, researchers have proposed the use of hybrid models that mix multiple architectures. These hybrids, such as the MambaFormer [32]—a mix of the popular SSM Mamba architecture with a standard Transformer—have shown potential in maintaining the desirable properties of multiple model classes.
@@
  • Failure to Use Pretrained Models. It is unclear how to integrate pretrained components from models with different architectures. Pretrained models are a key advantage of foundation models, but due to compatibility issues, hybrids are often trained from scratch, which is both limiting and costly.
A potential solution to the latter challenge is the use of model merging [45, 47, 44, 17, 8, 18], some of which can operate cross-architecture [1, 13]. Unfortunately, such tools are embryonic–they are expensive and it is unclear how well they work with the diverse architectures a user may seek to build a hybrid from.
We propose a framework for automatically designing hybrid architectures that overcomes these obstacles. Our approach is inspired by principles from neural architecture search (NAS), but applies these at the level of LM blocks rather than convolutional cells [26, 24] or operations [39, 38]. The resulting framework is simple, tractable, and it sidesteps merging different architectures by using simple projectors to translate between the “languages” spoken by various architectures. This enables us to include blocks from many different architectures/models with no changes required. Furthermore, inspired by the mechanistic architecture design framework (MAD) [37], we show how to learn hybrids via MAD that transfer to new tasks.
Concretely, with our proposed system, Manticore, we:
@@
    Automatically select language models, without training several models from scratch,
@@
    Automatically construct pretrained hybrids without evaluating the entire search space,
@@
    Explore when it is possible to program hybrids without full training.
Experimentally, our automatically designed hybrids compete with existing hybrids and models on the MAD tasks [37] and Long Range Arena [41, LRA], we produce pretrained hybrids that improve downstream fine-tuning performance on a variety of language tasks, and we show that Manticore can be programmed using MAD.
Figure 1: Manticore enables: (1) cross-architecture LM selection, (2) the construction of pretrained hybrids, and (3) the ability to program hybrids to have certain skills.
@@ 2 Related work
Language Model Architectures: Transfomers and Beyond. Transformers are currently the dominant LM architecture. The success of the “vanilla” architecture introduced by Vaswani et. al. [43] has led to many proposed variations. The quadratic complexity of the base self-attention operation has inspired the search for alternative architectures that offer comparable performance with subquadratic complexity. One line of work builds off state-space models, with variations made to enable language modeling [35, 36, 14, 3]. Another line of work involves linear-complexity attention by formulating transformers as RNNs and expressing self-attention as a kernel dot-product [20]. Other approaches increase the expressivity of this formulation with data-dependent gating [46]. Our work does not propose a new architecture. Instead, we focus on the idea that practitioners should be able to take advantage of new architectures in a transparent way.
Neural Architecture Search & Mechanistic Search. Neural architecture search (NAS) techniques are used to automatically search for optimal architectures. These techniques have produced state-of-the-art models in several different architectures and data domains. Much of the challenge in NAS is the complexity of the search procedures; in the most standard form, NAS involves a difficult bilevel optimization over a large search space. Much effort has been aimed at reducing these costs, often via continuous relaxations of the large search spaces, with efficient, end-to-end differentiable search techniques like DARTS [26], GAEA [24], and DASH [39].
Using NAS to discover architectures for language modeling—and especially those that may rival Transformers—has thus far been hard. A promising approach is the MAD suite [37], which uses “mechanistic tasks” (synthetic tasks organized around simple principles) to search for high-quality subquadratic architectures. While we do not seek to discover new architectures, we are inspired by this approach in our effort to search for hybrid architectures.
@@ 3 Methods
We now describe Manticore, our framework for automatically designing hybrid architectures by mixing components of pretrained models. We rely on projectors to align features across architectures, then apply a convex combination to aligned features, as shown in Figure 1.
In Section 3.1, we discuss and formally define the structure of Manticore hybrids: projectors and convex combination mixture weights, as well as how both components are used within Manticore. In Section 3.2, we detail the NAS-inspired search procedures and training routines involved in pretraining, fine-tuning, and programming hybrids. Finally, we provide the synthetic and real data settings that we use in our experiments in Section 4.
@@ 3.1 The Structure of Manticore Hybrids
Component Models We refer to a model that is used in Manticore as a component model. Any modern decoder-only LM can be used as a component model in our framework. In this section, we will formally define the general high-level structure of the component models that we support. For an LM $M$ with model embedding dimension $d_{M}$ on a sequence of $t$ tokens from a set $\mathcal{V}$, denoted $x=(x_{1},...,x_{t})\in\mathcal{V}^{t}$, a forward pass $M(x)$ is typically computed using the following recipe:
@@
This recipe applies to virtually all transformer-based LMs, recurrent models, and state-space models. Manticore supports all of these and any architecture that follows this recipe.
Projectors Suppose we have pretrained component models $M$ and $M^{\prime}$. Assuming that the model dimensions are the same for both models ($d_{M}=d_{M^{\prime}}$), blocks from $M$ and $M^{\prime}$ may not be compatible, as their input and output features are distributed differently. It is also possible that $d_{M}\neq d_{M^{\prime}}$, in which case composing blocks from $M$ and $M^{\prime}$ is not well defined.
To overcome this issue, we apply projectors to both the inputs and the outputs of a block (or a sequence of blocks, discussed in Section 3.1) that we wish to combine in Manticore hybrids. Overall, our goal in designing projectors is to enable the blocks of $M$ and $M^{\prime}$ to share a common representation, such that their features are compatible and can be reused in the resulting hybrid model. This is conceivably challenging—the mapping between feature spaces could be highly nonlinear and might require a lot of task-specific data to adequately learn the mapping. If the mapping is indeed highly nonlinear, we might need heavyweight multi-layer projectors with a large number of parameters. This could substantially increase parameter counts, inference cost, and could increase the data requirement for learning them. So do projectors need to be heavyweight, data-hungry, highly nonlinear objects? Fortunately, we find that the answer is no—we find that a simple linear transformation with a gated residual, pretrained on general language data, is sufficient. [^4] When this gating is combined with Equation 1, we see that the use of gated residuals ensures that the component architectures are still in our search space. This is a convenient property that allows Manticore to fall back on a component model when it outperforms hybrids.
@@
we want $\text{Proj-in}^{(\ell)}_{(k)}:\mathbb{R}^{t\times d_{\max}}\to\mathbb{R}^{t\times d_{M_{(k)}}}$ and $\text{Proj-out}^{(\ell+n)}_{(k)}:\mathbb{R}^{t\times d_{M_{(k)}}}\to\mathbb{R}^{t\times d_{\max}}$, so that
$\bigg(\text{Proj-out}^{(\ell+n)}_{(k)}\circ M_{(k)\text{Block}}^{(\ell+n)}\circ...\circ M_{(k)\text{Block}}^{(\ell)}\circ\text{Proj-in}^{(\ell)}_{(k)}\bigg):\mathbb{R}^{t\times d_{\text{max}}}\to\mathbb{R}^{t\times d_{\text{max}}}.$
@@
For input $x\in\mathbb{R}^{t\times d_{M_{(k)}}}$ we parameterize projectors as linear layers with gated residuals:
$\text{Proj-in}^{(\ell)}_{(k)}(x;\alpha):=$ $(1-\alpha)\cdot\text{Linear}_{d_{\text{max}}\to d_{M_{(k)}}}(x)+\alpha\cdot\text{Trunc}(x;d_{M_{(k)}})$
$\text{Proj-out}^{(\ell)}_{(k)}(x;\alpha):=$ $(1-\alpha)\cdot\text{Linear}_{d_{M_{(k)}}\to d_{\text{max}}}(x)+\alpha\cdot\text{Pad}(x;d_{\text{max}}).$
@@
Respectively, $\text{Trunc}(\cdot;d)$ and $\text{Pad}(\cdot;d)$ truncate and zero-pad input to dimension $d$, and $\text{Linear}_{d\to d^{\prime}}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ is a learnable linear layer with gating weights $\alpha\in[0,1]$. In total, where $\alpha\in\Delta^{K-1}$ and $I_{k}$ is a length-$n_{k}$ vector of block indices from component model $k$, we define the output of the block sequence defined by $I_{k}$ as
$h_{k}(x;\alpha_{k},I_{k})=$ $\bigg(\text{Proj-out}^{(I_{k,n_{k}})}_{(k)}\circ M_{(k)\text{Block}}^{(I_{k,n_{k}})}\circ...\circ M_{(k)\text{Block}}^{(I_{k,1})}\circ\text{Proj-in}^{(I_{k,1})}_{(k)}\bigg)(x;\alpha_{k}).$
@@
Mixture Weights Next, we would like to mix the activations of different component models’ block sequences, in a way that allows us to learn how much influence the blocks from each component model will have on the overall hybrid model. Learning the amount of influence that each block sequence should have on the overall hybrid is critical—if certain blocks produce less helpful features, we need a way to down-weight them. Conversely, we want to use the best blocks in our hybrid as much as possible—we want to up-weight helpful blocks. Overall, a parameterization that allows us to learn these weights should lead to better hybrids. We do this by taking a convex combination of the projectors’ outputs: given the projected features $h_{k}(x;\alpha_{k},I_{k})$ for each component model $k\in[K]$, we output a convex combination of projected features
$\text{Mix}_{\alpha}(x;I_{1},...,I_{K})=\sum_{k\in[K]}\alpha_{k}h_{k}(x;\alpha_{k},I_{k}).$ (1)
@@
We reuse the convex combination weights as the gating weights in the projectors. This choice yields the convenient property that when the mixture weights $\alpha$ are set to one in index $k$ and zero everywhere else, the Mix function exactly computes a sequence of blocks from component model $k$ while completely ignoring the projectors and the blocks from other component models. We adopt a popular parameterization for mixture weights from the NAS literature [26]: we parameterize $\alpha$ as a softmax of a parameter vector—that is, $\alpha_{k}:=\frac{\exp(a_{k})}{\sum_{j\in[K]}\exp(a_{j})}$ for all $k\in[K]$.
@@
Manticore We are now ready to define our overall hybrid architecture. We seek to create a hybrid from $K$ component models, $M_{(1)},...,M_{(K)}$, each with a potentially different number of blocks, denoted $L_{M_{(k)}}$ for component model $k$. We fix $L$ to be the number of Manticore blocks, where $L$ is a common factor of each of the depths $L_{M_{(k)}}$, for all $k\in[K]$—we treat this choice of factor as a hyperparameter. For each of the $L$ Manticore blocks, we want to mix a sequence of blocks from each of the $K$ component models. We also want the number of blocks from each model $k\in[K]$ that are allocated to a single Manticore block to be evenly spread throughout the $L$ Manticore blocks—this is why we require $L$ to be a factor of $L_{M_{(k)}}$. For each component model $k\in[K]$, divide the indices of the blocks $L_{M_{(k)}}]$ evenly into $L$ contiguous parts, denoted as $L_{M_{(k)}}]=(I_{k,1},...,I_{k,L})$. Then, adopting the notation from our component models, a Manticore block is defined as